How many 4-letter words can be formed from the letters A,B,C,D,E,F if the first letter must\nnot be a vowel and repetition of letters is not allowed
1
Expert's answer
2020-11-19T17:06:00-0500

There are 4 possible letters at the first place: B, C, D, F. Then

"4\cdot\dfrac{5!}{(5-3)!}=4(5)(4)(3)=240"

There 240 possible 4-letter words.